The percent of fat calories that a person consumes each day is normally distributed with a mean of about 33 and a standard deviation of about ten. Suppose that 9 individuals are randomly chosen. Let X = average percent of fat calories. 


(a) Give the distribution of X. 

(b) For the group of 9, find the probability that the average percent of fat calories consumed is more than four. 

Graph the situation and shade in the area to be determined. 

(c) Find the first quartile for the average percent of fat calories. 
Answer #1

Solution:

Ans:

Given that

mean  \mu = 33

standard deviation \sigma = 10

n = 9

a)sampling distribution of sample mean follow normal distribution with mean equal to population mean,\mu and standard deviation equal to \sigma / \sqrt{n} .

\mu_{\bar{x}}=33

\sigma_{\bar{x}}=10/\sqrt{9}=3.33

b)

z = (5 - 33) / 3.33 = -8.408

P(z>-8) = 1.0000

c)

For first quartile:

P( Z<=z ) = 0.25

z = normsinv (0.25) = -0.6745

P 25 = 33 - 0.6745 * 10 / sqrt (9)

P 25 = 30.75

> that's super hard to read

Jacob Greer Thu, Nov 4, 2021 9:22 PM